OpenStudy (anonymous):

Can someone check this please.... Does it make sense, work, or is the solution to : find the ratio of a square inscribed in an equilateral triangle. Area of a square = s² Area of a triangle = (t²√3)/4 30◦-60◦-90◦ rule T = s+2(s/(√3 ) ) Lets say s = 3 (since we are looking for a ratio and all sides are equal, we can use any number in the equation). T=3+2(3/(√3 ) ) = 6.46 Area of triangle = (6.46² x √3)/4=18 Area of square = 3² = 9 The ratio of the area of an equilateral triangle to the area of the inscribed square is 18/(9 )= 2/1=2:1

OpenStudy (anonymous):

I came up with this: Sin (60) = s/l l = s/Sin60 = (2s√3)(1/3) Similar triangles t/s = l + s/s = 2/3s√3/s = 2√3/3 +1 t = (2√3/3 + 1)s t2 = (2/√3 + 1)2 = (4/3 + 1 + 4/√3)s2 = (7/3 + 4√3/3)s2 Area of a triangle: A = t2√3/4 = (7/3 + 4√3/3)s2/4 *√3 = (7√3/12 + 1)s2 Area of a square s2/(7√3/12 + 1)s2 = 1/7√3 + 12/12 = (12/7√3 + 12)(7√3 – 12/7√3 – 12) Area of a triangle = 84√3 – 144/147 – 144 = 84√3 – 144/3 = 28√3-48 = 12 + 7√3/12

OpenStudy (anonymous):

Well Daniel helped me. Its a team project, and now I need to choose which is best.

myininaya (myininaya):

so what info is given to us? every thing before you talked about the area of a square? and you are showing two answers right?

OpenStudy (anonymous):

yes I am showing two different answers... and the question is... hang on... let me get it... A square with sides of length s is inscribed in an equilateral triangle with sides of length t. Find the exact ratio of the area of the equilateral triangle to the area of the square.

OpenStudy (anonymous):

Key word .... exact ratio?? The other team member came up with a 2:1 and I came up with a square root over a number... I'm just looking for a second opinion.

myininaya (myininaya):

k i'm almost to a conclusion please wait

OpenStudy (anonymous):

good business

myininaya (myininaya):

see if that helps

myininaya (myininaya):

so i would say the answer is (12+7sqrt{3}):12

OpenStudy (anonymous):

you came up with a Different answer?? oh no

OpenStudy (anonymous):

ahhhh good! ok ... did you understand the ratio part of my first solution... did it make sense at all?

OpenStudy (anonymous):

why would the lady say 2:1 ?

myininaya (myininaya):

what lady

OpenStudy (anonymous):

as her answer

OpenStudy (anonymous):

like i said its a team project the first answer was hers and the second was mine

OpenStudy (anonymous):

you were the tie breaker :)

myininaya (myininaya):

so the ratio i got is what you got?

myininaya (myininaya):

is that bottom thing your ratio the 12 is under everything not just that last part right?

OpenStudy (anonymous):

yeah with a square root kind of answer

OpenStudy (anonymous):

yes

myininaya (myininaya):

in the middle i wrote my ratio if you wanted it

myininaya (myininaya):

or if you didn't understand my conclusion without me writing my conlcusion there is my conclusion lol

OpenStudy (anonymous):

lol, thank you... I was right!!

myininaya (myininaya):

your way is different from mine but the conclusion we got is all that matters

OpenStudy (anonymous):

exactly... so what kind of resasoning do you have behind the first solution being ratio 2:1

OpenStudy (anonymous):

I need some math words to convince her my solution is right...

myininaya (myininaya):

ok so you want me to look and find fault with her logic

OpenStudy (anonymous):

if you could... as proof for me... please

myininaya (myininaya):

strike 1: it looks as if she rounded she has like 6.49 .....it says to find exact ....

myininaya (myininaya):

strike 2: when finding area of traingle she squared the base

OpenStudy (anonymous):

damn your good!

myininaya (myininaya):

the formula is 1/2*b*h

OpenStudy (anonymous):

yep

myininaya (myininaya):

you know if she didn't make those two mistakes she would had been good

OpenStudy (anonymous):

thank you.... I appreciate your time and help! :)

myininaya (myininaya):

actually i can understand her way alittle better because it was easier to compare her stuff to my notes because she had some similar expressions

OpenStudy (anonymous):

ok... very interesting...

myininaya (myininaya):

mean theres nothing wrong with your way lol

OpenStudy (anonymous):

no no... lol

myininaya (myininaya):

could you understood what i did

OpenStudy (anonymous):

yes... math on a deep level!

myininaya (myininaya):

:)

OpenStudy (anonymous):

Thank you again!

myininaya (myininaya):

your welcome

OpenStudy (amistre64):

didja get it ;)

OpenStudy (anonymous):

yes i sure did! How are you!

OpenStudy (amistre64):

Im fine thnx :) got some monies finally and now my car needs some work tho. The steering links have finally deteriorated to the point they need to be replaced. Need to get new track bar, tie rods, and a few bushings to get it back in shape.

OpenStudy (anonymous):

yeah that is no fun. I dropped 900 on my car... the control board went out and two front bushings! I could have found so much more to do with 900 buckaroos!! :)

OpenStudy (anonymous):

I am getting ready to go to Cancun! I will see you back in 5 days! Friday! I have one last week left in math! You will see me again!! Have a great week! Thanks for all your help!

OpenStudy (amistre64):

:) Good luck, and have fun ;)

OpenStudy (anonymous):

Hey do you have facebook?

OpenStudy (amistre64):

i do, but rarely use it lately ... Ynohtna Attertsim

OpenStudy (amistre64):

I use it mainly to spy on my kids lol

OpenStudy (amistre64):

I was thinking about making an account for just the math stuff tho; not sure how that would work...

OpenStudy (anonymous):

lol thats funny! The math thing would be neat!